OpenStudy (anonymous):

what is the equation of the line that passes through the points (-2,-3) and (4,-5 )

OpenStudy (anonymous):

In point slope, or slope intercept?

OpenStudy (ajprincess):

First find the slope using the formula \[m=\frac{y_2-y_1}{x_2-x_1}\] here \((x_1,y_1)\)=(-2,-3) and \((x_2,y_2)\)=(4,-5). Then choose one of the points. If u choose \((x_1,y_1)\) then equation can be found using \(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):